# LUCID: LLM-Generated Utterances for Complex and Interesting Dialogues

### v1

@@ Abstract
Virtual assistants are poised to take a dramatic leap forward in terms of their dialogue capabilities, spurred by recent advances in transformer-based Large Language Models (LLMs). Yet a major bottleneck to achieving genuinely transformative task-oriented dialogue capabilities remains the scarcity of high quality and linguistically sophisticated data. Existing datasets, while impressive in scale, have limited domain coverage and contain few genuinely challenging conversational phenomena; those which are present are typically unlabelled, making it difficult to assess the strengths and weaknesses of models without time-consuming and costly human evaluation. Moreover, creating high quality dialogue data has until now required considerable human input, limiting both the scale of these datasets and the ability to rapidly bootstrap data for a new target domain. We aim to overcome these issues with LUCID, a modularised and highly automated LLM-driven data generation system that produces realistic, diverse and challenging dialogues. We use LUCID to generate a seed dataset of 4,277 multi-domain, multi-intent conversations across 100 intents to demonstrate its capabilities. The generated conversations include a wide range of challenging phenomena and diverse user behaviour, identifiable via a set of turn-level tags. Finally, we provide separate test sets for seen and unseen intents, allowing for convenient out-of-distribution evaluation. We release both the data generation code and the dataset itself [^1] Please see http://github.com/apple/ml-lucid-datagen for the data and project code.
@@
[…]
@@ 1 Introduction
As AI virtual assistants become more sophisticated, there will be an increasing need for dialogue datasets with more challenging conversational phenomena that can be used for model fine-tuning and evaluation. Existing datasets include multi-turn, multi-intent and multi-domain conversations Rastogi et al. (2020); Budzianowski et al. (2018), in addition to multi-lingual datasets Goel et al. (2023); FitzGerald et al. (2023); Hung et al. (2022); Li et al. (2021). However, in each case, the number of intents covered is relatively small. Moreover, the conversational phenomena included in these datasets are often limited in scope. The PRESTO dataset (Goel et al., 2023) does explicitly label specific conversational phenomena, but to the best of our knowledge it is unique in this respect. Additionally, current machine-to-machine data collection methods still involve varying degrees of human involvement, with humans paraphrasing machine generated templates into natural language, and/or manually crafting plausible sequences of intents as dialogue outlines (Shah et al., 2018; Rastogi et al., 2020). Even this reduced level of human involvement in the data generation process, relative to even more onerous human-to-human or human-to-machine approaches, is costly and time consuming, and difficult to scale to new target domains.
@@
To overcome these issues, we introduce LUCID, LLM-generated Utterances for Complex and Interesting Dialogues. LUCID is composed of a pipeline of modularised LLM calls that create realistic, accurate and complex data, allowing the data generation process to scale to more intents, slots and challenging conversational phenomena. In particular, based on a simple description of an intent (for example ‘send an email’), LUCID automatically generates a schema for the intent, including a list of required and optional slot names, types and values. A mock back-end [^2] The mock back-end converts intents into Python classes, which are then instantiated as objects whose interface is exposed to the system agent is then created for the intent that interacts with LLM-driven system and user agents. The system agent produces system commands, which are then executed by the mock back-end.
To introduce diversity and complexity into the data generation process, conversation rules are provided to the user agent to guide it towards specific conversational phenomena. Adding new conversational phenomena to the system is a simple matter of adding a set of imperatives written in natural language to a python file, for example "When Lucid asks about {arg1}, don’t answer, but instead say something to another person who is in the same room" (see Figure 1).
@@
Ensuring data quality is a central challenge for a machine-to-machine generation process. LUCID addresses this issue by breaking down the generation process into a pipeline of multiple, simpler LLMs calls, thereby compartmentalising the data generation task into manageable steps that an LLM can consistently perform accurately. In addition, we use multiple LLM-based validators which discard conversations that might contain an issue. Our if in doubt, discard philosophy ensures a high quality standard for the data being created. Finally, we use sampling throughout our generation process to ensure variety in the length and content of each conversation. To showcase the quality, realism and sophistication of the data generated by LUCID, we release a seed dataset of 92,699 turns.
@@
To summarise our contributions: 1) We introduce a highly automated methodology for producing high quality LLM-generated conversations. 2) We provide a challenging dataset of 4,277 dialogues across 100 intents, including more labelled conversational phenomena than previous dialogue datasets. 3) We provide two tests sets, a test set for seen intents, and an additional test set for unseen intents, allowing for convenient out-of-distribution evaluation. 4) We release the data generation code to enable large scale data generation across different intents and domains, with the option of adding additional, complex conversational flows.
@@ 2.1 Task Oriented Dialogue Datasets
The most popular approach for creating dialogue datasets involves human-to-human interactions, with user annotators interacting with Wizard of Oz (WoZ) annotators Budzianowski et al. (2018); El Asri et al. (2017); Zhu et al. (2020); Eric et al. (2017); Wen et al. (2017). While using human annotators can create diverse, large scale datasets, this is done at a considerable cost, with expert annotators required for accurate dialogue annotations. User annotators follow a generated conversation plan (Budzianowski et al., 2018; El Asri et al., 2017; Zhu et al., 2020), guiding their interactions with the WoZ agent. These instructions can include how the user should navigate specific scenarios, for example when no results match their constraints (El Asri et al., 2017). We find that even in a purely machine-to-machine setup, generating conversation plans for each dialogue remains an effective way to ensure conversational variety. LUCID extends these plans to also specify how and when any complex conversational phenomena should be included in the dialogue.
@@ 2.2 Automated Data Collection Methods
Conversations can also be created based on user-machine interactions, for example with users telephoning computerized agents Henderson et al. (2014a, b). Recently, Goel et al. (2023) collected conversations between annotators and a virtual assistant simulator to generate natural language conversation histories, which provide context to semantically annotated user turns.
Human involvement in automated data generation includes ensuring the quality of the dataset, paraphrasing user and agent responses (Shah et al., 2018; Rastogi et al., 2020), providing semantic annotations (Goel et al., 2023; Budzianowski et al., 2018), outlining the sequences of user intents (Rastogi et al., 2020), and identifying out of scope or incoherent examples (Goel et al., 2023). We show that, with recent advances in language modelling OpenAI (2023), by reducing the data generation task into manageable steps, and using our if in doubt, discard validation methodology, it is now possible to achieve the same quality in an almost entirely machine-to-machine generation process.
2.3 Data Generation with LLMs
Wu et al. (2023) recently introduce a framework allowing the interaction of multiple, different LLMs, based on the idea that LLMs can solve highly challenging tasks if these tasks are broken into smaller steps. While Wu et al. (2023) are successful in generating dialogues for a group chat scenario, this does not require the intent and slot labelling needed for task-oriented dialogue. For other NLP tasks, to avoid labelling errors or poor quality data, generating data with LLMs can involve human annotators reviewing the generated utterances Liu et al. (2022); Wiegreffe et al. (2022), or using the generated data as unlabelled data to be used with knowledge distillation Meng et al. (2022); He et al. (2022); Stacey and Rei (2023). Labeled data generation has been successful for other tasks without human input Honovich et al. (2023); Ye et al. (2022); Schick and Schütze (2021); West et al. (2022); Wu et al. (2022), however noise may be an issue for a large proportion of the data Honovich et al. (2023); Schick and Schütze (2021).
@@ 3 Method
LUCID decomposes the data generation process into 14 individual LLM calls [^3] At this time we are unable to provide the details of the LLMs we used at each stage. If we receive this approval, we will provide an updated version of the paper., described here as stages, creating manageable steps that LLMs can perform accurately. Alongside our validation process, including our if in doubt, discard approach, reducing the complexity of each LLM call helps to ensure the quality of our generated data. The data generation process consists of four main components (see Figure 2): the generation of intents (stages 1-2), a conversational planner (stages 3-8), turn-by-turn generation of conversations (stages 9-12), and our validation process (stages 13 and 14). The turn-by-turn data generation involves a User LLM agent interacting with a System LLM agent, which in turn interacts with a mock back-end created for each of our generated intents.
@@ 3.1 Intent Generation
Schema for each intent are generated by an LLM, using a short human-authored natural language description of the intent (stage 1). In total, we provided LUCID with 54 transactional intent descriptions across 13 domains (see Appendix D). Using these descriptions, LUCID calls an LLM to generate the intent and slot names, as well as the data type of each slot and whether it is mandatory or optional. Each transactional intent has a corresponding query intent [^4] There are some cases when transactional intents have the same query intent - see Appendix D for details, with the purpose of retrieving entities created by that intent.
@@ 3.2 Conversation Planner
The conversation planner provides instructions that guide a user LLM agent down certain types of conversational flows. The planner specifies: 1) the sequence of intents, 2) the slot values for each intent, and 3) any complex conversational phenomena that must be included, specifying when and how these phenomena should be incorporated. This creates a plan that the user simulator must adhere to, reducing the complexity of the data generation task, while also ensuring variety in the generated conversations. This plan is communicated to the User LLM at each turn through a series of conversation rules, specifying the next intent (and its corresponding slot values) that should be requested, in addition to specifying the next complex conversational phenomena. The planner creates the conversation plan at the start of each conversation, involving a pipeline of different stages (stages 3-8).
3.2.1 Generating Conversation Structures
Firstly, the planner decides the sequence of intents that will be included in a conversation. Depending on the primary intent used to start the conversation, the planner then decides which intents are likely to follow this intent, with the aim of creating both varied and realistic conversations.
To achieve this, the planner is provided with a sample of intents across all domains. An LLM then generates a plausible sequence of intents from this sample (stage 3). To ensure variety in the generated conversations, the planner makes further use of sampling, choosing how many intents should be provided in the conversation, which optional slots should be discussed, the conversational phenomena (both happy and unhappy paths) that should be included, and which slots and intents any conversational phenomena should be applied to.
Figure 3: Examples for eight of the nine challenging conversational phenomena included in the LUCID dataset. We also included ‘cancellation’ examples which are similar to ‘delay confirmation’, resulting in the system not confirming a given intent.
@@ 3.2.2 Generating Slot Values
The slot values chosen by the planner substantially impact the conversations, and as a result, we have multiple, separate stages for generating the slot values (stages 4, 5, 6 & 7). Choosing slot values that are not plausible or coherent leads to unrealistic conversations. Moreover, to ensure variety in our generated conversations, we require variety within the slot values decided by the planner.
This process involves updating the slot values to make sure these are realistic and coherent (stage 4), generating a reason why the user wants to perform any subsequent intents (stage 5), and generating slot values for the subsequent intents based on this justification (stage 6). Finally, an LLM updates the slot values across every intent in the conversation to ensure they are consistent and realistic when considered collectively (stage 7). See Appendix E for further details.
We use an LLM to generate realistic entities to be returned after any queries. These entities will be provided to the Response LLM to communicate back to the user (stage 8).
@@ 3.3 Generating Conversations
Conversations are generated turn-by-turn with a User LLM interacting with a System LLM, which in turn interacts (via Pythonic function calls and variable assignments) with a mock back-end for each intent, and a Response LLM communicating natural language responses back to the user. The user behaviour is governed by the conversational rules created by our conversation planner, shaping the outcome of the conversations.
Conversations start with an utterance from the User LLM (stage 9), which is then interpreted and labelled by the System LLM (stage 10). Initially, no string slot values are predicted. These values are predicted in an additional stage (stage 11), where an LLM is instructed, where possible, to choose the string values from spans of the user utterance, providing fewer opportunities for the LLM to hallucinate. The predicted semantic labels then interact with a mock back-end for each intent, assessing whether any additional slots are required or whether the user should instead be asked to provide confirmation. The mock back-end provides these instructions to a Response LLM (stage 12), which generates a natural language response to the user.
@@ 3.4 LLM-based Validation
We implement an LLM-based validation process to ensure reliable and consistent labelling in the conversations, using our principle of if in doubt, discard. First, based on the observation that the system LLM is more uncertain about incorrect predictions, we repeat the system predictions twice (using temperature values less than 1), and abort the conversation if the three predictions are not identical (stage 13). Additional validation is then performed by another LLM (stage 14) which also labels the user requests, except this time with access to the conversation rules that the user is following. Predictions must exactly match the original System LLM outputs, otherwise the conversation is discarded.
Further validation is provided for our conversational phenomena, described in Section 3.5, in addition to a range of post-processing rules introduced after a qualitative analysis (described in Section 4.3). In total, 56% of the conversations generated passed our validation checks (see Appendix H for further information).
To avoid wasting valuable conversational data, we salvage the prefix of an aborted conversation up to the point where the validation error was identified [^5] To avoid overly short conversations, we do this only if at least one intent has been performed already or at least 10 turns have occurred. In these cases, we truncate the conversation, sampling LLM generated natural language responses that justify interrupting the conversation.
@@ 3.5 Introducing Additional Conversational Phenomena
To make interesting, diverse and challenging conversations, we introduce a wide range of conversational phenomena which are labelled automatically at the turn level (see Figure 3). These phenomena include sarcastic or irrelevant replies, or cases where the user is overheard in another conversation. LUCID also contains examples where a user corrects themselves, either within a turn (in-turn correction) or in a later turn (correction). Alternatively, a user may cancel an intent (cancellation) or delay confirming the intent until a future turn (delay confirmation). Our conversational phenomena also consider how the user responds to prompts from the virtual assistant. If the virtual assistant requests a value for one slot, the user can respond about a different slot (respond different slot). See Table 3 for the full distribution of these phenomena in our data.
The conversation rules are sampled by the planner for each conversation, before the planner assigns these phenomena to specific intents (and slots) within the conversation. In addition to the unhappy path conversational phenomena discussed, we also introduce different happy paths via conversation rules to introduce variety within conversations. These rules state whether or not the user should provide some, all or no slot values when expressing the intent for the first time.
We introduce additional validations to ensure these challenging turns are correctly predicted by our LLMs. When the conversation rules instruct a user to introduce a specific phenomenon for a certain slot value, the user is instructed to also provide a signal (in the form of a special token) to show that this unhappy path is being performed. We then use this signal for validation purposes, ensuring that the following system command matches the expected response for this phenomenon. However, we do not provide these special tokens to the system which interprets and labels the user request; these would not be available to a real virtual assistant and we find that including them during data generation can result in unrealistic target labels (e.g. if a user’s ‘irrelevant’ response accidentally constitutes a plausible slot value).
Figure 4: An example labeled conversation (a simplified example from the dataset). Each dialogue contains user turns, system, signal and response turns.
@@ 3.6 Annotation Scheme and our Mock Back-end
We apply a concise labelling system to track the states for each intent in a conversation. This labelling system follows a Pythonic syntax, with function calls used to initialise intents and entities when these are first mentioned, and attribute assignments used for any subsequent slot filling operations (see Figure 4). This schema-based, concise form of semantic labelling is highly convenient, and avoids the need for state tracking for each individual turn. The Pythonic variables also provide a natural way to encode anaphoric phenomena such as pronominal coreference, as well as effectively leveraging the LLM’s pretraining data, in which Python is likely very well represented.
Our labeling involves four types of turns in a conversation: 1) User turns; 2) System turns, labelling the intentions of a user (including slot values and the intents specified); 3) Signal turns, returned after our mock back-end processes the system command, recommending the next action; and 4) Response turns, which are natural language responses back to the user.
LUCID automatically creates the mock back-end for each intent using the schema generated in steps 1 and 2. This involves generating a Python class to represent the intent in question, which is then instantiated as an object and interacts with the system commands to indicate when mandatory slots have not yet been provided, or when confirmation is still required before the intent can be performed.

### v2

@@ Abstract
Spurred by recent advances in Large Language Models (LLMs), virtual assistants are poised to take a leap forward in terms of their dialogue capabilities. Yet a major bottleneck to achieving genuinely transformative task-oriented dialogue capabilities remains the scarcity of high quality data. Existing datasets, while impressive in scale, have limited domain coverage and contain few genuinely challenging conversational phenomena; those which are present are typically unlabelled, making it difficult to assess the strengths and weaknesses of models without time-consuming and costly human evaluation. Moreover, creating high quality dialogue data has until now required considerable human input, limiting both the scale of these datasets and the ability to rapidly bootstrap data for a new target domain. We aim to overcome these issues with LUCID, a modularised and highly automated LLM-driven data generation system that produces realistic, diverse and challenging dialogues. We use LUCID to generate a seed dataset of 4,277 conversations across 100 intents to demonstrate its capabilities, with a human review finding consistently high quality labels in the generated data [^1] Please see http://github.com/apple/ml-lucid-datagen for the data and project code.
@@
Joe Stacey1thanks: Work undertaken while author was an intern at Apple         Jianpeng Cheng2        John Torr2        Tristan Guigue2 Joris Driesen2        Alexandru Coca3∗        Mark Gaynor2        Anders Johannsen2 1Imperial College London      2Apple            3University of Cambridge j.stacey20@imperial.ac.uk, ac2123@cam.ac.uk {jianpeng.cheng,jtorr,tguigue}@apple.com {joris_driesen,mgaynor,ajohannsen}@apple.com
@@ 1 Introduction
As AI virtual assistants become more sophisticated, there is an increasing need for dialogue datasets with more challenging conversational phenomena for both fine-tuning and evaluation. Existing datasets include multi-turn, multi-intent and multi-domain conversations Rastogi et al. (2020); Budzianowski et al. (2018), in addition to multi-lingual datasets Goel et al. (2023); FitzGerald et al. (2023); Hung et al. (2022); Li et al. (2021). However, in each case, the number of intents covered is relatively small. Moreover, the conversational phenomena included in these datasets are often limited in scope [^2] The PRESTO dataset (Goel et al., 2023) does explicitly label specific conversational phenomena, but to the best of our knowledge it is unique in this respect. Additionally, current machine-to-machine data collection methods still involve varying degrees of human involvement, with humans paraphrasing machine generated templates into natural language, and/or manually crafting plausible sequences of intents as dialogue outlines (Shah et al., 2018; Rastogi et al., 2020).
@@
To overcome these issues, we introduce LUCID, LLM-generated Utterances for Complex and Interesting Dialogues. LUCID is composed of a pipeline of modularised LLM calls that create realistic, accurate and complex data, allowing the data generation process to scale to more intents, slots and challenging conversational phenomena. LUCID involves automated intent generation, with a mock back-end [^3] The mock back-end converts intents into Python classes, which are then instantiated as objects created for each intent. This mock back-end then interacts with LLM-based user and system agents, generating dialogues without the need for human annotation.
@@
Ensuring data quality is a central challenge for a machine-to-machine generation process. We address this issue by breaking down the generation process into a pipeline of multiple, simpler LLM calls, thereby compartmentalising the data generation task into manageable steps that an LLM can consistently perform accurately. In addition, we use multiple LLM-based validators which discard conversations that might contain an issue. Our if in doubt, discard philosophy ensures a high quality standard for the data being created.
@@
We release the data generation code to enable large scale data generation across different intents and domains, with the option of adding additional, complex conversational flows. We also provide training data, validation data, and two tests sets, a test set for seen intents, and an additional test set for unseen intents, allowing for convenient out-of-distribution evaluation.
@@ 2.1 Task Oriented Dialogue Datasets
The most popular approach for creating dialogue datasets involves human-to-human interactions, with user annotators interacting with Wizard of Oz (WoZ) annotators Budzianowski et al. (2018); El Asri et al. (2017); Zhu et al. (2020); Eric et al. (2017); Wen et al. (2017). While using human annotators can create diverse, large scale datasets, this is done at a considerable cost, with expert annotators required for accurate dialogue annotations. User annotators follow a generated conversation plan (Budzianowski et al., 2018; El Asri et al., 2017; Zhu et al., 2020), guiding their interactions with the WoZ agent. We find that even in a purely machine-to-machine setup, generating conversation plans for each dialogue remains an effective way to ensure conversational variety.
@@ 2.2 Automated Data Collection Methods
Human involvement in automated data generation includes ensuring the quality of the dataset, paraphrasing user and agent responses (Shah et al., 2018; Rastogi et al., 2020), providing semantic annotations (Goel et al., 2023; Budzianowski et al., 2018), outlining the sequences of user intents (Rastogi et al., 2020), and identifying out of scope or incoherent examples (Goel et al., 2023). We show that, with recent advances in language modelling OpenAI (2023); Ouyang et al. (2022), by reducing the data generation task into manageable steps, and using our if in doubt, discard validation methodology, it is now possible to achieve the same quality in an almost entirely machine-to-machine generation process. Parallel work by Liu et al. (2024) also introduces an automated method for generating task-oriented dialogue data. While we focus on the accurate labelling of challenging and diverse conversations, Liu et al. (2024) consider a variety of user personas with different styles of system responses.
See Appendix A for related work generating data with LLMs for other tasks.
@@ 3 Method
LUCID decomposes the data generation process into 14 individual LLM calls, described here as stages, creating manageable steps that LLMs can perform accurately. Alongside our if in doubt, discard validation, reducing the complexity of each LLM call helps to ensure the quality of our generated data. The data generation process consists of four main components (see Figure 2): the generation of intents (stages 1-2), a conversational planner (stages 3-8), turn-by-turn generation of conversations (stages 9-12), and our validation process (stages 13 and 14). The turn-by-turn data generation involves a User LLM agent interacting with a System LLM agent, which in turn interacts with a mock back-end created for each intent.
@@ 3.1 Intent Generation
Schema for each intent are generated by an LLM, using a short human-authored natural language description of the intent (stage 1) [^4] Our code also allows intent schema to be created manually. Using these descriptions, LUCID calls an LLM to generate the intent and slot names, as well as the data type of each slot and whether it is mandatory or optional. In total, 100 intents are generated across 13 domains (see Appendix H for a detailed breakdown of how transactional and query intents are generated).
@@ 3.2 Conversation Planner
The conversation planner provides instructions that guide a user LLM agent down certain types of conversational flows. The planner specifies: 1) the sequence of intents, 2) the slot values for each intent, and 3) any complex conversational phenomena that must be included, specifying when and how these phenomena should be incorporated. This creates a plan that the user must adhere to, reducing the complexity of the data generation task, while also ensuring variety in the generated conversations. This plan is communicated to the User LLM at each turn through a series of conversation rules.
The planner also decides the sequence of intents that will be included in a conversation (stage 3). Depending on the primary intent used to start the conversation, the planner then decides which intents are likely to follow this intent, with the aim of creating both varied and realistic conversations. See Appendix B for further details about the planner.
@@ 3.2.1 Generating Slot Values
The slot values chosen by the planner substantially impact the conversations, and as a result, we have multiple, separate stages for generating the slot values (stages 4, 5, 6 & 7). This process involves updating the slot values to make sure these are realistic and coherent (stage 4), generating a reason why the user wants to perform any subsequent intents (stage 5), and generating slot values for the subsequent intents based on this justification (stage 6). Finally, an LLM updates the slot values across every intent in the conversation to ensure they are consistent and realistic when considered collectively (stage 7).
We additionally use an LLM to generate realistic entities to be returned after any queries (stage 8).
@@ 3.3 Generating Conversations
Conversations are generated turn-by-turn with a User LLM interacting with a System LLM, which in turn interacts (via Pythonic function calls and variable assignments) with a mock back-end for each intent. A Response LLM then communicates natural language responses back to the user. The user behaviour is governed by the conversational rules created by our planner, shaping the outcome of the conversations.
Conversations start with an utterance from the User LLM (stage 9), which is then interpreted and labelled by the System LLM (stage 10). Initially, no string slot values are predicted. These values are predicted in an additional stage (stage 11), where an LLM is instructed, where possible, to choose the string values from spans of the user utterance (avoiding hallucinations). The predicted semantic labels then interact with a mock back-end for each intent. The mock back-end then informs the System LLM about any missing slots or whether confirmation is required for the intent. Finally, the Response LLM responds back to the user (stage 12), requesting any additional slots or asking the user for confirmation.
@@ 3.4 LLM-based Validation
We implement an LLM-based validation process to ensure reliable and consistent labelling in the conversations, using our principle of if in doubt, discard. First, based on the observation that the system LLM is more uncertain about incorrect predictions, we repeat the system predictions twice (using a temperature value of 0.7), and abort the conversation if the three predictions are not identical (stage 13). Additional validation is then performed by another LLM (stage 14) which also labels the user requests, except this time with access to the conversation rules that the user is following. These predictions must also exactly match the original System LLM predictions, otherwise the conversation is aborted. The validation in stages 13 and 14 is performed before the string slot values are predicted, avoiding conversations being aborted when these slot values have slightly different phrasing. Further validation is described in Appendix E.
@@ 3.5 Introducing Additional Conversational Phenomena
To make interesting, diverse and challenging conversations, we introduce a wide range of conversational phenomena which are labelled automatically at a turn level (see Figure 4). These phenomena include sarcastic or irrelevant replies, or cases where the user is overheard in another conversation. LUCID also contains examples where a user corrects themselves, either within a turn (in-turn correction) or in a later turn (correction). Alternatively, a user may cancel an intent (cancellation) or delay confirming the intent until a future turn (delay confirmation). Our conversational phenomena also include cases where the virtual assistant requests a value for one slot, but the user responds about a different slot (respond different slot). Finally we also include ASR early end errors (ASR-early end), where the LLM produces truncated slot values where the user text is abruptly cut off. See Table 4 for the full distribution of these phenomena.
Figure 3: A (simplified) example labelled conversation. Each dialogue contains user, system, signal and response turns.
@@ 3.6 Annotation Scheme and our Mock Back-end
We apply a concise labelling system to track the states for each intent in a conversation. This labelling system follows a Pythonic syntax, with function calls used to initialise intents and entities when these are first mentioned, and attribute assignments used for any subsequent slot filling operations (see Figure 3). This schema-based, concise form of semantic labelling is highly convenient, and avoids the need for state tracking for each individual turn.
Our labelling involves four types of turns: 1) User turns; 2) System turns, labelling user intentions; 3) Signal turns, returned after our mock back-end processes the system command; and 4) Response turns, which are natural language responses back to the user. Further details on our labelling schema is provided in Appendix F.